### v1

# Cross-View UAV Geo-Localization with Precision-Focused Efficient Design: A Hierarchical Distillation Approach with Multi-view Refinement

Abstract
Cross-view geo-localization (CVGL) enables UAV localization by matching aerial images to geo-tagged satellite databases, which is critical for autonomous navigation in GNSS-denied environments. However, existing methods rely on resource-intensive fine-grained feature extraction and alignment, where multiple branches and modules significantly increase inference costs, limiting their deployment on edge devices. We propose Precision-Focused Efficient Design (PFED), a resource-efficient framework combining hierarchical knowledge transfer and multi-view representation refinement. This innovative method comprises two key components: 1) During training, Hierarchical Distillation paradigm for fast and accurate CVGL (HD-CVGL), coupled with Uncertainty-Aware Prediction Alignment (UAPA) to distill essential information and mitigate the data imbalance without incurring additional inference overhead. 2) During inference, an efficient Multi-view Refinement Module (MRM) leverages mutual information to filter redundant samples and effectively utilize the multi-view data. Extensive experiments show that PFED achieves state-of-the-art performance in both accuracy and efficiency, reaching 97.15% Recall@1 on University-1652 while being over more efficient in FLOPs and faster than previous top methods. Furthermore, PFED runs at 251.5 FPS on the AGX Orin edge device, demonstrating its practical viability for real-time UAV applications. The project is available at https://github.com/SkyEyeLoc/PFED.
I Introduction
Cross-view geo-localization (CVGL) aims to determine the geographic location of a query image by matching it against a geo-tagged reference database. For drones, this capability is especially critical, offering a pathway to autonomous localization where Global Navigation Satellite System (GNSS) signals are unavailable or unreliable. The task typically involves matching multi-view drone-level images to a corresponding top-down satellite image, a process complicated by extreme variations in viewpoint, scale, and appearance.
To improve the cross-view matching precision, the field has rapidly evolved from early methods using handcrafted descriptors to dominant deep learning paradigms built on Siamese or Triplet networks [1, 2, 3, 4, 5]. More recently, Vision Transformers (ViTs) [6] and their variants, such as TransGeo [7] and FSRA [8], have set high performance standards by leveraging self-attention to learn powerful, view-invariant global representations. However, despite these advancements, several critical challenges remain that hinder the deployment of these models in practical, real-world scenarios.
Firstly, the pursuit of higher accuracy has led to an escalating computational burden. Many state-of-the-art methods achieve superior performance by incorporating complex auxiliary branches, resource-demanding cross-view alignment strategies. As shown in Figure 1, while models like CCR [9] and Mean [4] achieve high recall, they do so with immense computational overhead (e.g., over 90G FLOPs). This paradigm is fundamentally ill-suited for resource-constrained edge platforms like drones, where memory, power, and latency are primary concerns, creating a significant gap between algorithmic advancement and practical applicability.
Secondly, existing methods [10, 11] often lack an explicit mechanism to address the inherent trade-off between semantic abstraction and spatial fidelity. As features pass through deeper layers, they gain semantic robustness at the cost of losing the spatial details (e.g., rooftop textures, landmark patterns) that are critical for discriminating between visually similar locations. Furthermore, the significant data imbalance and domain discrepancy between the different views lead to asymmetric convergence, resulting in suboptimal feature alignment.
Thirdly, current approaches often make inefficient use of multi-view information. While a sequence of drone images provides a rich representation of a landmark, most methods either process each view independently [12, 13] or resort to computationally prohibitive techniques like 3D reconstruction to fuse views [14]. The former fails to leverage the collaborative potential of multiple perspectives, while the latter imposes a computational barrier that is impractical for real-time onboard applications. There is a clear need for a lightweight mechanism that can intelligently select and aggregate the most informative views without significant processing overhead.
Based on the above analysis, a straightforward idea is to simply deploy a lightweight network, but this approach typically leads to a significant performance collapse. To bridge this performance gap, we propose the Precision-Focused Efficient Design (PFED), a novel framework that fundamentally redefines the accuracy-efficiency trade-off by shifting complexity from the inference stage to the training stage. As illustrated in Figure 2, our approach comprises two key innovations. 1) During training, we introduce Hierarchical Distillation for Cross-View Geo-Localization (HD-CVGL). This paradigm first employs Fine-Grained Inverse Self-distillation (FISD) to compel the model’s final, semantically rich layer to retain the critical spatial details discovered by its shallower layers. Concurrently, our Uncertainty-Aware Prediction Alignment (UAPA) mechanism adaptively balances the learning between drone and satellite branches based on predictive uncertainty, fostering a more stable and robust training process. Finally, we perform cross-distillation from a domain-specialized foundation model to distill generalized knowledge into our lightweight student network, boosting its precision without incurring any additional inference overhead. 2) During inference, we propose a lightweight Multi-view Refinement Module (MRM). This post-processing method leverages mutual information to intelligently select a small, diverse, and highly informative subset of drone views. By aggregating the features from only these key views, MRM significantly enhances localization accuracy while drastically reducing the feature matching cost, enabling efficient and precise multi-view utilization.
Overall, this paper makes the following contributions:
(1) We propose the Precision-Focused Efficient Design (PFED) paradigm, which resolves the accuracy-efficiency dilemma in CVGL by concentrating model complexity in the training phase, resulting in a highly accurate yet computationally lightweight inference model.
(2) We introduce Hierarchical Distillation for CVGL (HD-CVGL), a novel training framework that synergistically combines inverse self-distillation, uncertainty-aware alignment, and cross-distillation to create a compact feature extractor that excels at capturing both semantic and spatial information.
(3) We propose the Multi-view Refinement Module (MRM), an efficient post-processing method, and theoretically demonstrate based on manual information how it enhances localization by optimally selecting and fusing multi-view drone information while minimizing feature matching overhead.
(4) The empirical results on the University-1652 and SUES-200 benchmarks demonstrate that PFED establishes a new state-of-the-art in both accuracy and efficiency. Additional deployment tests on edge devices further validate PFED’s real-time performance capabilities.
II Related work
II-A Cross-view UAV Geo-localization
Cross-View Geo-Localization involves matching a query image against a geo-referenced database from another viewpoint, a foundational challenge within the field of computer vision. Early methods relied on handcrafted local descriptors like SIFT [15] or global scene descriptors like HoG [16], but showed limited robustness to the extreme cross-view domain gap. The paradigm shifted with deep metric learning, where Siamese or Triplet networks [17, 18] learn a shared, view-invariant embedding space. To explicitly address geometric inconsistencies, some works integrate geometric priors, notably using a polar transform to align satellite and ground-level perspectives [19, 20] before feature comparison. While ground-to-satellite matching [21, 22] sets the foundation, the unique oblique perspective of Unmanned Aerial Vehicles (UAVs) introduces a more severe and spatially non-uniform multi-view set of geometric challenges.
UAV-to-Satellite Geo-Localization. This task is particularly challenging due to the multi-view oblique and low-altitude perspective of UAVs, which creates significant domain gaps between platforms. Following the establishment of key benchmarks [1, 12], research has progressed from improving feature robustness with attention mechanisms [23] to the now-dominant Vision Transformer (ViT) architectures. ViTs like TransGeo [7] and FSRA [24] leverage self-attention to learn powerful global representations, setting high performance standards. More recently, the field has explored more efficient and powerful backbone architectures. For instance, several works have successfully employed ConvNeXt [25] to extract highly discriminative global features, achieving a strong balance of performance and efficiency [26]. This trend highlights a focus on backbone potency for feature extraction. However, most of contemporary methods [4, 27, 13] achieve higher accuracy by incorporating sophisticated auxiliary branches or modules, this paradigm is ill-suited for resource-constrained platforms. We follow a Precision-Focused Efficient Design (PFED) paradigm to empower a simple backbone. By strengthening its feature extraction capabilities without adding any inference-time modules, our model achieves competitive precision while drastically reducing the computational burden, making it ideal for deployment on platforms like UAVs.
II-B Multi-view Refinement
Traditional UAV-based geo-localization relies on direct matching between individual query images and a reference database [28, 3, 8]. However, these approaches struggle with significant viewpoint discrepancies caused by factors such as occlusions from structures or vegetation, and the diverse perspectives captured by UAVs operating at varying altitudes and angles, making accurate location recognition increasingly challenging. Recent works have recognized the importance of leveraging multiple drone views to improve localization accuracy with approaches using 3D reconstruction [29, 30, 31] to represent scenes from multi-view observations and iteratively refining camera poses to align rendered views with satellite imagery. This multi-view fusion paradigm [32] demonstrates significant improvements in localization performance by exploiting the rich geometric and appearance information contained across different viewpoints.
However, despite their theoretical advantages, these multi-view approaches impose substantial computational burdens that severely limit their practical UAV onboard applicability. In contrast to these complex approaches with substantial computational barriers, we propose an efficient multi-view refinement module (MRM) that operates as a lightweight post-processing method. Rather than constructing expensive 3D representations or performing iterative rendering optimization, our method directly aggregates selected features from multiple drone viewpoints through Mutual Information (MI).
Mutual information has emerged as a principled criterion for selecting informative views in a wide variety of complex systems. In 3D reconstruction, mutual information has guided next-best-view selection effectively. Recent work proposes gaussian splatting shannon mutual information for real-time assessment of visual mutual information from novel viewpoints, enabling principled next-best-view selection without prior information [33]. For contrastive learning, Tian et al. [34] argue that we should reduce the mutual information between views while keeping task-relevant information intact, proposing the InfoMin principle. Similarly, in multi-view clustering, recent work [35] minimizes MI between common and view-specific representations to exploit inter-view complementary information while maximizing MI between refined and original representations to preserve principal information.
Existing UAV-based geo-localization methods either process all available views through expensive reconstruction-based methods [14, 29], missing the opportunity for principled information-theoretic view selection. In this work, we bridge this gap by proposing an efficient mutual information-based post-processing method that jointly optimizes view selection and feature aggregation for UAV-based geo-localization.
III Proposed Method
In this section, we provide a comprehensive introduction to our proposed PFED. As illustrated in Figure 2, we follow a precision-focused efficient design in two stages: training and inference post-processing. During the training stage, we propose Hierarchical Distillation for CVGL (HD-CVGL). Initially, we employ Fine-Grained Inverse Self-distillation (FISD) based on the internal stages of an efficient student model. Concurrently, to address the challenge of data imbalance, we introduce Uncertainty-Aware Prediction Alignment (UAPA), achieving a more robust and stable train. Subsequently, guided by a fine-tuned foundation model, we perform cross-distillation to obtain a student model with higher precision. In the inference stage, we propose a Multi-view Refinement Module (MRM) to better leverage multi-view information and further reduce the time cost of feature matching.
III-A Hierarchical Distillation for CVGL
Fine-Grained Inverse Self-distillation
In the CVGL task, a fundamental trade-off in designing deep networks exists between semantic abstraction and spatial fidelity. As shown in Figure 2 (c), the student network , composed of hierarchical stages (), transforms an input image into a sequence of feature representations , where . As the depth increases, gains semantic abstraction at the cost of losing the fine-grained spatial details present in the shallower features. In cross-view matching, these discarded low-level details often contain critical view-invariant cues (e.g., rooftop textures, landmark patterns).
To ensure the final representation is both semantically robust and perceptually detailed, we propose a novel method named fine-grained inverse self-distillation (FISD), a form of hierarchical consistency regularization. This approach inverts the conventional knowledge transfer paradigm [36]. Instead of the deep layer teaching the shallow, we compel the final student layer to retain the discriminative knowledge discovered by the shallower ”teacher” layers. This inverse knowledge transfer is motivated by two fundamental observations:
Targeting the Task-Specific Layer: It is the final layer’s feature that is ultimately used for the matching task. Consequently, this is the representation that must be refined and enriched to maximize performance.
Leveraging a Spatial-Detail-Preserving Teacher: Shallower layers serve as an authoritative teacher by preserving the fine-grained spatial information that deeper, more semantic layers progressively lose to abstraction.
We attach an auxiliary classification head to each stage’s feature map , producing logits . The core of FISD is to align the probability distribution of the final stage, , with those from all preceding stages .
Hybrid Loss Function. The training objectives of HD-CVGL in the first stage (FISD) are threefold, integrating multi-stage supervision with a focused deep metric learning objective on the final feature representation, as follows:
Multi-Level Supervision and Self-Distillation. To ensure that features at all levels of the hierarchy are semantically meaningful, we first apply a standard Cross-Entropy (CE) loss to the logits from each of the stages. This deep supervision loss is a weighted sum:
Further, to regularize the network, we employ intra-model inverse self-distillation. We first define the softened probability distribution for any stage using a temperature :
Then minimizes the Kullback-Leibler (KL) divergence between teacher distribution and the student distribution :
where denotes the KL divergence and are hyper-parameters that weight the contribution of each shallow teacher. This loss regularizes the learning of the final layer, forcing it to produce predictions consistent with the detail-oriented perspectives of the early stages.
Refining Embeddings with Symmetric Metric Learning. While the supervisory losses guide class predictions, we introduce a dedicated metric learning objective, , applied exclusively to the final-stage embeddings. The goal is to learn a shared, domain-invariant embedding space where features from drone-view and satellite-view images of the same landmark are aligned. Let a batch consist of pairs of geographically corresponding images, , where is a drone-view image and is its satellite-view counterpart. Let denote the final-stage feature embedding.
First, to enforce intra-domain class separability, we apply the triplet loss with hard sample mining strategy [11, 1]. This loss operates within each domain (drone and satellite) to ensure that embeddings of the same class are clustered together, while embeddings of different classes are pushed apart. It is applied to the final features from both domains:
Second, to achieve cross-view alignment, we employ a Symmetric InfoNCE Loss [37], a standard and powerful objective in CVGL. For a drone feature anchor , its corresponding satellite feature serves as the positive sample. All other satellite features in the batch, , act as negatives. The loss is computed symmetrically, using satellite features as anchors as well. For simplicity, let and . The Cross-view Symmetric Contrastive (CSC) loss is composed of two symmetric terms: a drone-to-satellite loss () and a satellite-to-drone loss (). The overall loss is their average:
Each directional loss is formulated as an InfoNCE loss. For a batch of pairs, the drone-to-satellite loss is defined as:
and the satellite-to-drone loss is its symmetric counterpart:
where is the cosine similarity and is a temperature parameter. This loss explicitly pulls the representations of corresponding drone-satellite pairs together while pushing them away from non-matching pairs. The total deep metric learning objective combines these components, applied to the final embeddings from both views:
This composite loss ensures that the final embedding space is both class-discriminative within each domain and robustly aligned across different views.
Uncertainty-Aware Prediction Alignment
In cross-view geo-localization, a significant challenge arises from the inherent imbalance and domain discrepancy between drone and satellite views. Specifically, for each drone-view query, only a single positive satellite sample exists within a large gallery, creating a severe data imbalance. Conventional methods often treat both domains equally, overlooking the fact that one view might be inherently more ambiguous than the other, leading to suboptimal feature alignment. As shown in Figure 3, our observation suggests an asymmetric convergence behavior, where the model may be specializing on the features of the dominant domain. This insight motivates balanced feature learning for robust cross-view matching. To address this, we introduce an uncertainty-aware prediction alignment mechanism that adaptively modulates the learning process based on the model’s confidence in each view.
Our approach begins by quantifying the predictive uncertainty of each domain. Inspire by [38, 36, 39], we employ Shannon entropy, a standard measure of uncertainty, calculated from the softmax probabilities derived from the model’s output logits. For a given logit vector over classes, the uncertainty is defined as:
here, represents the predicted probability for class , and is the corresponding logit. A higher entropy value signifies greater uncertainty and thus lower confidence in the prediction.
We then dynamically adjust the alignment process based on the relative uncertainty between the drone and satellite branches. We compute the uncertainties and for the respective logit predictions and . The core of our method is cross-modal self distillation with an adaptive temperature scaling strategy. The temperature is adjusted based on the uncertainty gap, :
where is a pre-defined base temperature and is the sigmoid function. The sigmoid function smoothly maps the unbounded uncertainty gap to a bounded scaling factor in the range (0, 1). This formulation increases the temperature when the drone-view model is more uncertain than the satellite-view model (i.e., ). This is critical for our self-distillation. When the drone view is ambiguous (e.g., due to view change or occlusions), its model is naturally uncertain. Forcing it to match the satellite’s high-confidence prediction would create a conflicting learning signal. Raising the temperature softens the target, mitigating this conflict and providing a more appropriate guidance for the uncertain student.
Finally, we use the adaptive temperature to guide the alignment between the two domains via a Kullback-Leibler (KL) divergence loss. The satellite branch acts as a ”teacher,” providing a soft target distribution for the drone ”student” branch. The alignment loss, , is formulated as:
By making the alignment process sensitive to predictive uncertainty, our method fosters a more robust and stable training process, effectively mitigating the challenges posed by data imbalance and domain-specific ambiguity.
Cross-distillation training
The second step of our hierarchical framework is Cross-Distillation Training (CDT), a process designed to transfer knowledge from a large foundation model (DINOv2-base [40]) teacher to a lightweight ConvNeXt-tiny [25] student, as illustrated in Figure 2 (a). This approach adheres to the PFED paradigm by directly training the final student architecture without requiring any inference-time modifications. Critically, this teacher is not used off-the-shelf; it is first specialized for the target task through a parameter-efficient fine-tuning process on the University-1652 dataset, where only the final two Transformer blocks were made trainable. This approach creates an expert teacher that retains vast, general visual knowledge while acquiring high-level semantic understanding specific to our domain.
To ensure comprehensive knowledge transfer, we distill information at both the feature and logit levels as follows:
where denotes normalization, denote the teacher’s and student’s final stage output feature, are their respective temperature-scaled probability outputs, and is the ground-truth label. By distilling both feature semantics and classification logic from the final outputs of a domain-adapted teacher, our student model learns highly relevant representations and decision boundaries for the task.
III-B Multi-view Refinement Module (MRM)
In drone-based visual geo-localization, capturing multiple viewpoints of landmarks is essential for robust matching against satellite references. During data collection, drones systematically capture images at predetermined positions, resulting in a comprehensive multi-view representation. Let denote the set of aerial views captured at different heights and azimuth angles . While this dense sampling ensures complete coverage of landmarks, processing all views during inference poses significant computational challenges, particularly for real-time navigation applications.
To address this challenge, as shown in Figure 4, we propose the Multi-view Refinement Module (MRM), a post-processing technique that intelligently selects an optimal subset with . The key insight is that not all views contribute equally to geo-localization accuracy: some perspectives capture more distinctive features, while others may be less informative due to occlusions or viewing angles. We formulate this view selection problem within an information-theoretic framework, maximizing the mutual information between landmark identities while ensuring spatial diversity.
Theoretical Foundation. Our approach is grounded in information theory [41, 42], where the goal is to select views that maximize the mutual information between view features and landmark labels :
where is the differential entropy of view features and is the conditional entropy given the landmark class. Direct computation of mutual information for high-dimensional features is computationally prohibitive [43]. We propose an efficient approximation based on the theoretical connection between Fisher discriminant ratio and mutual information under structured assumptions.
Proposition 1. Under the assumption that view features follow class-conditional Gaussian distributions with equal covariance, the mutual information can be lower-bounded by:
where and are the between-class and within-class variances for view .
This theoretical insight enables us to use computationally efficient statistics as proxies for mutual information while maintaining theoretical rigor.
Information-Theoretic View Importance Assessment. Given extracted multi-view features for a landmark, the MRM quantifies each view’s information content through three complementary measures grounded in information theory. First, we approximate the marginal entropy of view features through log-variance:
This measure captures the information richness of the view, with higher entropy indicating more diverse visual patterns.
Second, we estimate the dynamic entropy through the log-range of feature activations:
This complements the variance-based entropy by capturing the span of feature activations, identifying views with strong, distinctive features.
Most importantly, we compute the Gaussian-MI approximation to measure geo-discriminability:
where the between-class variance quantifies separation between different landmarks:
and the within-class variance measures consistency within each landmark class:
here, denotes the number of samples in class , is the mean feature for class in view , is the global mean, and is the total number of samples. This formulation provides a computationally efficient estimate of mutual information while maintaining theoretical guarantees under Gaussian assumptions.
The final information score integrates these measures, where denotes min-max normalization:
Submodular Optimization for Spatial Diversity. While information scores identify informative views, optimal subset selection must balance information content with spatial coverage. We formulate this as a submodular optimization problem that jointly maximizes information and diversity.
We model each view’s spatial position as , where represents altitude and the azimuth angle. The spatial distance between views incorporates both vertical and angular separation:
where accounts for circular angles, with weights , .
Proposition 2. The objective function
is submodular, and the greedy algorithm achieves a -approximation to the optimal subset.
The greedy selection iteratively adds views that maximize the marginal gain:
with balancing information content and spatial diversity.
Information-Weighted Multi-view Aggregation. After selecting the optimal subset , we perform information-weighted aggregation that reflects each view’s contribution to the mutual information. The aggregation weights are computed using a softmax function over information scores:
where is a temperature parameter controlling the sharpness of the weighting. This exponential weighting amplifies the contribution of high-information views while maintaining differentiability. The refined representation for landmark is:
Theoretical Analysis and Guarantees. Our manual information based framework provides the following theoretical guarantees:
Approximation Quality: Under Gaussian assumptions, the approximation error is bounded by where measures deviation from Gaussianity.
Computational Efficiency: By reducing the feature matching complexity from to , our MRM achieves a significant speedup. is the number of drone views of one landmark, this corresponds to a reduction in computational cost, rendering the approach viable for real-time applications.
The effectiveness of our approach stems from a principled connection between mutual information and discriminative learning. This connection enables efficient view selection, preserving the most informative perspectives while ensuring comprehensive spatial coverage of drone-view landmarks.
IV Experiment
IV-A Implementation Details
Datasets. We conduct extensive experiments on two prominent drone-based geo-localization benchmarks that offer complementary characteristics for comprehensive evaluation. University-1652 [1] is the first drone-based geo-localization dataset and SUES-200 [12] represents a pioneering benchmark that considers aerial photography captured by drones at different flight heights. These datasets collectively enable comprehensive evaluation of cross-view geo-localization algorithms under varying conditions—University-1652 excels in scale and multisource integration, while SUES-200 provides critical insights into altitude robustness and real-world applicability.
Experimental Setup. We train our model using a batch size of , where each batch contains different location IDs with samples per ID. This results in drone images and satellite images per batch. All images are resized to pixels for both training and testing phases. The model is trained for epochs using the SGD optimizer, initialized with a learning rate of and incorporating a 5 epoch warm-up phase following [44] to stabilize early-stage gradient dynamics. For cross-region evaluation in the unsupervised domain adaptation setting, we implement a contrastive learning fine-tuning strategy, limiting adaptation to a single epoch on the University-1652 dataset to mitigate distribution-specific overfitting. We employ the AutoAugment [45] strategy, which applies optimized combinations of augmentation operations including random cropping, translation, horizontal flipping, and color jittering. All experiments are conducted using the PyTorch framework on a single NVIDIA GeForce RTX 3090 GPU with 24GB memory.
Evaluation Protocol. Following standard practice in cross-view geo-localization [1, 7], we evaluate both retrieval accuracy and computational efficiency across all models
Retrieval Metrics. We adopt Recall@1 to measure the fraction of queries where the correct result appears in the top position. Average Precision (AP) provides a comprehensive assessment by considering all relevant items in the ranking.
Efficiency Metrics. We report three efficiency metrics: (1) Parameters: the total number of learnable parameters in millions; (2) FLOPs: floating-point operations computed for a single forward pass; (3) Processing Speed (FPS): in the feature extraction context, the number of UAV and satellite images processed per second. Note that we set the batch size to 16 to accommodate the multi-view UAV data processing.
| Method | Drone Satellite | Satellite Drone | |||||
| Recall@1 | AP | Recall@1 | AP | ||||
| LPN [2] [TCSVT’2021] | 62.39 M | 65.39 G | 218 | 77.71 | 80.80 | 90.30 | 78.78 |
| FSRA [8] [TCSVT’2021] | 53.16 M | 98.05 G | 100 | 85.50 | 87.53 | 89.73 | 84.94 |
| MCCG [28] [TCSVT’2023] | 56.65 M | 51.04 G | 313 | 89.40 | 91.07 | 95.01 | 89.93 |
| MuSe-Net [46] [PR’2024] | 82.90 M | 42.37 G | 405 | 74.48 | 77.83 | 88.02 | 75.10 |
| SCPNet [47] [TCSVT’2025] | - | - | - | 79.96 | 83.04 | 87.33 | 79.87 |
| TriSA [3] [TCSVT’2024] | 51.13 M | 43.18 G | 275 | 90.08 | 91.56 | 96.01 | 90.12 |
| Safe-Net [48] [TIP’2024] | 52.67 M | 24.58 G | 282 | 86.98 | 88.85 | 91.22 | 86.06 |
| SDPL [10] [TCSVT’2024] | 42.56 M | 69.71 G | 519 | 90.16 | 91.64 | 93.58 | 89.45 |
| SRLN [49] [TGRS’2024] | 193.03 M | - | - | 92.70 | 93.77 | 95.14 | 91.97 |
| Sample4Geo [13] [ICCV’2023] | 87.57 M | 90.24 G | 144 | 92.65 | 93.81 | 95.14 | 91.39 |
| CCR [9] [TCSVT’2024] | 156.57 M | 160.61 G | - | 92.54 | 93.78 | 95.15 | 91.80 |
| DAC [27] [TCSVT’2024] | 96.50 M | 90.24 G | 128 | 94.67 | 95.50 | 96.43 | 93.79 |
| MEAN [4] [TGRS’2025] | 36.50 M | 26.18 G | 307 | 93.55 | 94.53 | 96.01 | 92.08 |
| PFED (Ours) | 28.57 M | 4.45 G | 1022 | 93.87 | 94.83 | 95.72 | 92.57 |
| PFED (Ours w/ Post-process) | 28.57 M | 4.45 G | 1022 | 97.15 | 97.50 | 95.58 | 96.27 |
IV-B Comparison with State-of-the-Art Methods
We conduct a comprehensive evaluation of our proposed Precision-Focused Efficient Distillation (PFED) framework against a wide array of recent state-of-the-art methods. As presented in Table I, the comparison focuses on both retrieval accuracy (Recall@1, AP) and model efficiency (Parameters, FLOPs, FPS). Our results demonstrate that PFED not only establishes a new state-of-the-art in accuracy but also fundamentally redefines the efficiency frontier for this task.
Superior Efficiency. Our core model, PFED, exhibits high computational efficiency. With only 28.57M parameters and an exceptionally low 4.45G FLOPs, it is by far the most lightweight and computationally inexpensive model among all compared methods. To put this in perspective, compared to the recent efficient model MEAN [4], PFED reduces FLOPs by a remarkable factor of 5.8 and parameters by 21.2%. This optimization directly translates to a massive 3.3 increase in FPS, reaching 1022 images/second, a critical capability for real-time deployment on resource-constrained platforms.
State-of-the-Art Accuracy. Building upon this highly efficient foundation, our full model, PFED†, incorporates the Multi-view Refinement Module (MRM) as a post-processing step achieves an impressive 97.15% Recall@1 and 97.50% AP in the primary DroneSatellite retrieval task. This represents a substantial absolute improvement of 3.30% in R@1 over our efficient baseline and significantly surpasses the previous best-performing method, DAC [27], all while operating with over 20 fewer FLOPs (4.45G vs. 90.24G).
In summary, the results in Table I unequivocally show that PFED successfully achieves a trade-off between accuracy and efficiency in cross-view geo-localization, making it a highly practical and effective solution for real-world applications.
IV-C Unsupervised Domain Adaptation Results
To rigorously assess the generalization capabilities of our model, we conducted zero-shot evaluations by training on the University-1652 dataset and directly testing on the SUES-200 dataset without any fine-tuning. This challenging setting simulates real-world deployment where models must handle unseen data domains. The comparative results are detailed in Table II (DroneSatellite) and Table III (SatelliteDrone).
As demonstrated in Table II, our proposed PFED exhibits exceptional generalization in the DroneSatellite task. While MEAN [4] shows strong performance at 150m, our model consistently outperforms all state-of-the-art methods at 200m, 250m, and 300m. Notably, at the 300m range, PFED achieves a top-ranking R@1 score of 97.87, surpassing the next-best competitor by a significant 3.24 percentage points. It is crucial to highlight that PFED utilizes only 4.45G FLOPs, compared to the 26.18G FLOPs required by MEAN. This demonstrates that our model’s feature representation is more robust to severe scale variations while being remarkably lightweight.
The superior generalization of PFED is more pronounced in the SatelliteDrone task, as shown in Table III. In this scenario, our model unequivocally achieves the best performance across all evaluation altitudes. For instance, it surpasses the strong baseline DAC [27] by 3.75 percentage points in R@1 at the 250m altitude. This state-of-the-art accuracy is delivered with a model that is substantially more efficient, requiring only 28.57M parameters compared to DAC’s 96.50M.
IV-D Multi-weather UAV imagery degradation Results
As shown in Figure 5, to evaluate the model’s robustness in practical scenarios, we conducted extensive experiments on the University-1652 dataset under various environmental degradations, including fog, rain, darkness, and wind. As presented in Table IV, these tests assess the model’s ability to maintain high accuracy even when the visual quality of UAV imagery is severely compromised.
In the DroneSatellite retrieval task, our proposed PFED demonstrates remarkable resilience and establishes a new state-of-the-art across all tested conditions. From normal weather to the most severe degradations like darkness and combined fog with rain, PFED consistently achieves the highest Recall@1 and AP scores. For instance, under dark conditions, it outperforms the next-best method, MEAN [4], by a substantial margin of 5.37 percentage points in R@1 (93.27% vs. 87.90%). This superior performance is even more impressive given its efficiency; it delivers these results while requiring 5.8 times fewer FLOPs than MEAN.
In the more challenging SatelliteDrone task, PFED continues to show highly competitive performance. In the most complex scenarios such as ”Fog+Rain” and ”Wind”, PFED takes the lead in both metrics, achieving the best R@1 and AP scores. This proves that even when satellite images must be matched against heavily distorted drone views, our model’s feature representation remains robust and effective.
| DroneSatellite | |||||||||||
| 150m | 200m | 250m | 300m | ||||||||
| Model | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP | |||
| MCCG [28] [TCSVT’2023] | 56.65 M | 51.04 G | 313 | 57.62 | 62.80 | 66.83 | 71.60 | 74.25 | 78.35 | 82.55 | 85.27 |
| Sample4Geo [13] [ICCV’2023] | 87.57 M | 90.24 G | 144 | 70.05 | 74.93 | 80.68 | 83.90 | 87.35 | 89.72 | 90.03 | 91.91 |
| DAC[27] [TCSVT’2024] | 96.50 M | 90.24 G | 128 | 76.65 | 80.56 | 86.45 | 89.00 | 92.95 | 94.18 | 94.53 | 95.45 |
| MEAN [4] [TGRS’2025] | 36.50 M | 26.18 G | 307 | 81.73 | 87.72 | 89.00 | 91.00 | 92.13 | 93.60 | 94.63 | 95.76 |
| PFED (Ours) | 28.57 M | 4.45 G | 1022 | 79.85 | 83.32 | 91.10 | 92.79 | 95.57 | 96.45 | 97.87 | 98.32 |
| SatelliteDrone | |||||||||||
| 150m | 200m | 250m | 300m | ||||||||
| Model | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP | |||
| MCCG [28] [TCSVT’2023] | 56.65 M | 51.04 G | 313 | 61.25 | 53.51 | 82.50 | 67.06 | 81.25 | 74.99 | 87.50 | 80.20 |
| Sample4Geo [13] [ICCV’2023] | 87.57 M | 90.24 G | 144 | 83.75 | 73.83 | 91.25 | 83.42 | 93.75 | 89.07 | 93.75 | 90.66 |
| DAC[27] [TCSVT’2024] | 96.50 M | 90.24 G | 128 | 87.50 | 79.87 | 96.25 | 88.98 | 95.00 | 92.81 | 96.25 | 94.00 |
| MEAN [4] [TGRS’2025] | 36.50 M | 26.18 G | 307 | 91.25 | 81.50 | 96.25 | 89.55 | 95.00 | 92.36 | 96.25 | 94.32 |
| PFED (Ours) | 28.57 M | 4.45 G | 1022 | 92.50 | 83.81 | 97.50 | 91.75 | 98.75 | 94.59 | 97.50 | 96.04 |
| Method | DroneSatellite | SatelliteDrone | |||
| FLOPs | Recall@1 | AP | Recall@1 | AP | |
| (a) Normal | |||||
| LPN [2] | 65.39 G | 74.33 | 77.60 | 87.02 | 75.19 |
| MuSeNet [46] | 42.37 G | 74.48 | 77.83 | 88.02 | 75.10 |
| Sample4Geo[13] | 90.24 G | 90.55 | 92.18 | 95.86 | 89.86 |
| MEAN[4] | 26.18 G | 90.81 | 92.32 | 96.58 | 89.93 |
| PFED (Ours) | 4.45 G | 93.87 | 94.83 | 95.72 | 92.57 |
| (b) Fog | |||||
| LPN [2] | 65.39 G | 69.31 | 72.95 | 86.16 | 71.34 |
| MuSeNet [46] | 42.37 G | 69.47 | 73.24 | 87.87 | 69.85 |
| Sample4Geo[13] | 90.24 G | 89.72 | 91.48 | 95.72 | 88.95 |
| MEAN[4] | 26.18 G | 90.97 | 92.52 | 96.00 | 89.49 |
| PFED (Ours) | 4.45 G | 92.95 | 94.08 | 95.72 | 91.17 |
| (c) Rain | |||||
| LPN [2] | 65.39 G | 67.96 | 71.72 | 83.88 | 69.49 |
| MuSeNet [46] | 42.37 G | 70.55 | 74.14 | 87.73 | 71.12 |
| Sample4Geo[13] | 90.24 G | 85.89 | 88.11 | 94.44 | 85.71 |
| MEAN[4] | 26.18 G | 88.19 | 90.05 | 95.15 | 88.87 |
| PFED (Ours) | 4.45 G | 91.26 | 92.61 | 94.58 | 87.22 |
| (d) Fog+Rain | |||||
| LPN [2] | 65.39 G | 64.51 | 68.52 | 84.59 | 66.28 |
| MuSeNet [46] | 42.37 G | 65.59 | 69.64 | 85.02 | 67.78 |
| Sample4Geo[13] | 90.24 G | 85.88 | 88.16 | 93.44 | 85.27 |
| MEAN[4] | 26.18 G | 86.75 | 88.84 | 93.58 | 86.91 |
| PFED (Ours) | 4.45 G | 90.68 | 92.11 | 94.86 | 85.67 |
| (f) Dark | |||||
| LPN [2] | 65.39 G | 53.68 | 58.10 | 82.88 | 52.05 |
| MuSeNet [46] | 42.37 G | 53.85 | 58.49 | 80.74 | 53.01 |
| Sample4Geo[13] | 90.24 G | 87.90 | 89.87 | 96.01 | 87.06 |
| MEAN[4] | 26.18 G | 87.90 | 89.87 | 96.29 | 89.87 |
| PFED (Ours) | 4.45 G | 93.27 | 94.34 | 95.44 | 89.95 |
IV-E Anti-offset Generalization Results
In real-world UAV-view geo-localization, the captured query image is often not perfectly centered or aligned with its corresponding satellite-view gallery image. This spatial misalignment can be caused by variations in camera angle or differences in viewpoint. A robust model must be able to generalize well despite such spatial offsets.
To rigorously evaluate our model’s robustness against this challenge, we adopt the evaluation protocol popularized by SDPL [10]. As shown in Figure 6, the mapping from a desired shift to the required (left, top, right, bottom) padding tuple is as follows: top-left , top-right , bottom-left , and bottom-right .
The comprehensive results are presented in Table V. The findings demonstrate the superior robustness of our proposed PFED model. For example, under the most severe shift of pixels, where other models experience a significant performance collapse, PFED maintains an exceptional Rank-1 accuracy of . This represents a massive improvement of +16.92% over the second-best method, SDPL.
It is crucial to note that while the performance of all models degrades as the offset magnitude increases, PFED’s degradation is significantly more graceful. This highlights its ability to learn features that are inherently invariant to spatial translations. This anti-offset generalization capability, coupled with its significantly lower computational cost (4.45 GFLOPs) compared to competitors like FSRA (98.05 GFLOPs) and SDPL (69.71 GFLOPs), underscores the practical value and efficiency of our method for real-world deployment.
| Padding Pixel | FSRA (98.05 G) | LPN (65.39 G) | SDPL (69.71 G) | PFED (4.45 G) | ||||
| Recall@1 | AP | Recall@1 | AP | Recall@1 | AP | Recall@1 | AP | |
| (0,0) | 86.41 | 88.34 | 77.99 | 80.84 | 85.19 | 87.43 | 93.87 (+7.46) | 94.83 (+6.49) |
| (+20,0) | 85.51 | 87.59 | 76.64 | 79.62 | 84.42 | 86.78 | 93.19 (+7.68) | 94.28 (+6.69) |
| (+40,0) | 82.77 | 85.30 | 72.94 | 76.36 | 82.46 | 85.13 | 90.84 (+8.07) | 92.34 (+7.04) |
| (+60,0) | 77.95 | 81.18 | 66.85 | 70.91 | 79.22 | 82.43 | 85.17 (+5.95) | 87.59 (+5.16) |
| (-20,-20) | 84.35 | 86.62 | 76.40 | 79.42 | 84.39 | 86.76 | 93.87 (+9.48) | 94.83(+8.07) |
| (-40,-40) | 78.10 | 81.24 | 70.27 | 74.03 | 81.75 | 84.55 | 93.88 (+12.13) | 94.84 (+10.29) |
| (-60,-60) | 67.73 | 71.97 | 59.56 | 64.34 | 76.94 | 80.46 | 93.86 (+16.92) | 94.82 (+14.36) |
| (+20,-20) | 84.23 | 86.55 | 76.34 | 79.37 | 84.32 | 86.72 | 93.19 (+8.87) | 94.28 (+7.56) |
| (+40,-40) | 77.90 | 81.09 | 70.36 | 74.10 | 81.62 | 84.46 | 90.85 (+9.62) | 92.35 (+8.31) |
| (+60,-60) | 67.29 | 71.62 | 59.61 | 64.42 | 76.80 | 80.38 | 85.16 (+8.36) | 87.58 (+7.20) |
| (-20,+20) | 83.46 | 85.85 | 74.74 | 77.92 | 82.95 | 85.49 | 92.72 (+9.26) | 93.88 (+8.03) |
| (-40,+40) | 74.47 | 78.05 | 65.13 | 69.32 | 77.00 | 80.46 | 86.62 (+9.62) | 88.77 (+8.31) |
| (-60,+60) | 58.05 | 63.27 | 50.19 | 55.56 | 66.87 | 71.71 | 72.36 (+5.49) | 76.25 (+4.54) |
| HD-CVGL | MRM | DroneSatellite | SatelliteDrone | ||||
| SD | UAFA | CD | Recall@1 | AP | Recall@1 | AP | |
| ✗ | ✗ | ✗ | ✗ | 86.44 | 88.69 | 93.72 | 85.13 |
| ✓ | ✗ | ✗ | ✗ | 91.46 | 92.91 | 94.57 | 90.49 |
| ✓ | ✓ | ✗ | ✗ | 91.93 | 93.29 | 95.14 | 91.35 |
| ✓ | ✓ | ✓ | ✗ | 93.87 | 94.83 | 95.72 | 92.57 |
| ✓ | ✓ | ✓ | ✓ | 97.15 | 97.50 | 95.58 | 96.27 |
IV-F Ablation Studies
Our PFED framework is architected training-time optimization and lightweight inference-time post-process. To validate the specific contributions of each phase, we conducted extensive ablation studies, with the results summarized in Table VI.
Effectiveness of the Training-Phase Framework. The core objective of our training strategy is to distill rich, view-invariant knowledge into a single, efficient network, avoiding the need for multi-branch architectures during inference. Our analysis starts with a baseline model (row 1) achieving 86.44% Recall@1. By introducing Self-Distillation (SD), performance improves significantly to 91.46%, which ensures that the model can fully leverage its own multi-level feature hierarchy for more robust representation learning. Following this, integrating the UAFA module further advances the Recall@1 to 91.93%. Finally, Cross-Distillation (CD) elevates performance to 93.87%, completing the HD-CVGL framework. This step-wise improvement validates our training-phase design, which produces a powerful feature extractor ready for efficient inference. As visualized in Figure 7 and Figure 8, our PFED model learns a more discriminative feature distribution compared to the baseline. The features produced by PFED exhibit smaller intra-class variance and larger inter-class distance, indicating a well-structured feature space.
Effectiveness of the Inference-Phase Module. After establishing a strong student model through our training strategy, we evaluate the contribution of the MRM. The final row of Table VI shows the impact of applying MRM to the descriptors generated by our fully trained model. The result is a remarkable jump in performance to 97.15% Recall@1. This +3.28% gain over the already powerful base model demonstrates that by refining the feature set at inference time, MRM significantly enhances localization precision, fulfilling the ”Precision-Focused” aspect of our PFED paradigm without altering the underlying network architecture. Figure 9 illustrates the performance of different selected views. The model achieves its peak performance with 40 views. Even at its lowest setting (20 views), our approach still significantly outperforms the baseline without the MRM module. As demonstrated by the top-5 retrieval results on the University-1652 and SUES-200 datasets (Figure 10 and Figure 11), PFED exhibits robust cross-view matching capabilities.
Edge Device Deployment. To demonstrate the practical applicability of our approach, as shown in Figure 12, we evaluate PFED on two representative edge platforms: NVIDIA Jetson TX2 and AGX Orin. As shown in Table VII, our method achieves exceptional efficiency with only 4.45 GFLOPs, representing a 95.1% reduction compared to DAC. This dramatic reduction translates directly to superior real-time performance: PFED achieves 27.1 FPS on the resource-constrained TX2 (7.7 faster than DAC’s 3.5 FPS) and an impressive 251.5 FPS on AGX Orin (6.1 faster than DAC). Even compared to the efficient MEAN method, our approach demonstrates 2.9 and 2.8 speedups on TX2 and Orin respectively.
IV-G Limitations and Future Works
Despite its promising performance, PFED has several limitations that guide our future work. The current model, while excelling at DroneSatellite task exhibits an asymmetric performance due to its focus on the multi-view refinement of drone features. Future efforts will explore the intra-relationships within the satellite data to build a more generalized cross-view feature space. Additionally, PFED’s lightweight design faces challenges with extreme low-altitude distortions, which could be mitigated by developing adaptive, scale-aware mechanisms or pre-transformation techniques. Finally, to bridge the gap between efficient image-level retrieval and the need for precise localization, our ultimate goal is to evolve beyond static images. We plan to leverage multiview sequence data and develop an end-to-end framework that integrates coarse scene identification with fine-grained pose estimation, enabling robust and accurate geo-localization within large-scale satellite imagery.
V Conclusion
In this paper, we introduced Precision-Focused Efficient Design (PFED), a framework that resolves the critical accuracy-efficiency trade-off in cross-view geo-localization for UAVs. Our core strategy shifts computational complexity to the training stage, enabling superior performance on resource-constrained devices. We achieve this through two innovations: 1) A comprehensive Hierarchical Distillation (HD-CVGL) strategy during training, which incorporates our Uncertainty-Aware Prediction Alignment (UAPA) to robustly handle data imbalance and domain discrepancies, producing a highly discriminative yet compact student network without any inference overhead. 2) A lightweight Multi-view Refinement Module (MRM) at inference, which uses mutual information to select and fuse the most informative views, boosting accuracy while minimizing feature matching cost. Experiments show PFED sets a new state-of-the-art, achieving 97.15% Recall@1 on University-1652 while being over more FLOPs-efficient and faster than previous methods. Crucially, its 251.5 FPS on an AGX Orin validates its practicality for real-time UAV-based geo-localization in GNSS-denied environments.