### v2

# MobileGeo: Exploring Hierarchical Knowledge Distillation for Resource-Efficient Cross-view Drone Geo-Localization

Abstract
Cross-view geo-localization (CVGL) enables drone localization by matching aerial images to geo-tagged satellite databases, which is critical for autonomous navigation in GNSS-denied environments. However, existing methods rely on resource-intensive feature alignment and multi-branch architectures, incurring high inference costs that limit their deployment on mobile edge devices. We propose MobileGeo, a mobile-friendly framework designed for efficient on-device CVGL. MobileGeo achieves its efficiency through two key components: 1) During training, a Hierarchical Distillation (HD-CVGL) paradigm, coupled with Uncertainty-Aware Prediction Alignment (UAPA), distills essential information into a compact model without incurring inference overhead. 2) During inference, an efficient Multi-view Selection Refinement Module (MSRM) leverages mutual information to filter redundant views and reduce computational load. Extensive experiments demonstrate that MobileGeo outperforms previous state-of-the-art methods, achieving a 4.19% improvement in AP on University-1652 dataset while being over 5 more efficient in FLOPs and 3 faster. Crucially, MobileGeo runs at 251.5 FPS on an NVIDIA AGX Orin edge device, demonstrating its practical viability for real-time on-device drone geo-localization. The code is available at https://github.com/SkyEyeLoc/MobileGeo.
I Introduction
Cross-view geo-localization (CVGL) aims to determine the geographic location of a query image by matching it against a geo-tagged reference database. For drones, this capability is especially critical, offering a pathway to autonomous localization where Global Navigation Satellite System (GNSS) signals are unavailable or unreliable. The task typically involves matching multi-view drone-level images to a corresponding top-down satellite image, a process complicated by extreme variations in viewpoint, scale, and appearance.
To improve the cross-view matching precision, the field has rapidly evolved from early methods using handcrafted descriptors to dominant deep learning paradigms built on Siamese or Triplet networks [1, 2, 3, 4, 5]. More recently, Vision Transformers (ViTs) [6] and their variants, such as TransGeo [7] and FSRA [8], have set high performance standards by leveraging self-attention to learn powerful, view-invariant global representations. However, despite these advancements, several critical challenges remain that hinder the deployment of these models in practical, real-world mobile scenarios.
Firstly, the pursuit of higher accuracy has led to an escalating computational and resource burden. In Figure 2, many state-of-the-art methods achieve superior performance by incorporating complex auxiliary branches, resource-demanding cross-view alignment strategies. As shown in Figure 1, while models like CCR [11] and Mean [4] achieve high recall, they do so with immense computational overhead (e.g., over 90G FLOPs). This paradigm is fundamentally ill-suited for resource-constrained edge platforms, such as drones, where memory, power, and latency are critical concerns, creating a significant gap between algorithmic advancements and their practical applicability in mobile computing.
Secondly, existing methods [9, 12] often lack an explicit mechanism to address the inherent trade-off between semantic abstraction and spatial fidelity. As features pass through deeper layers, they gain semantic robustness at the cost of losing the spatial details (e.g., rooftop textures, landmark patterns) that are critical for discriminating between visually similar locations. Furthermore, the significant data imbalance and domain discrepancy between the different views lead to asymmetric convergence, resulting in suboptimal feature alignment.
Thirdly, current approaches often make inefficient use of multi-view information. While a sequence of drone images provides a rich representation of a landmark, most methods either process each view independently [13, 14] or resort to computationally prohibitive techniques like 3D reconstruction to fuse views [15]. The former fails to leverage the collaborative potential of multiple perspectives, while the latter imposes a computational barrier that is impractical for real-time onboard applications. There is a clear need for a lightweight mechanism that can intelligently select and aggregate the most informative views without significant processing overhead.
Based on the above analysis, a straightforward idea is to simply deploy a lightweight network, but this approach typically leads to a significant performance collapse. To bridge this performance gap, we propose the MobileGeo, a novel mobile-friendly framework that fundamentally redefines the accuracy-efficiency trade-off by shifting complexity from the inference stage to the training stage. As illustrated in Figure 3, our approach comprises two key innovations. 1) During training, we introduce Hierarchical Distillation for Cross-View Geo-Localization (HD-CVGL). This paradigm first employs Fine-Grained Inverse Self-Distillation (FISD) to compel the model’s final, semantically rich layer to retain the critical spatial details discovered by its shallower layers. Concurrently, our Uncertainty-Aware Prediction Alignment (UAPA) mechanism adaptively balances the learning between drone and satellite branches based on predictive uncertainty, fostering a more stable and robust training process. Finally, we perform cross-distillation from a domain-specialized foundation model to distill generalized knowledge into our lightweight student network, boosting its precision without incurring any additional inference overhead. 2) During inference, we propose a lightweight Multi-view Selection Refinement Module (MSRM). This post-processing method leverages mutual information to intelligently select a small, diverse, and highly informative subset of drone views. By aggregating the key features, MSRM significantly enhances accuracy while drastically reducing the feature matching cost, enabling mobile-efficient multi-view utilization.
Overall, this paper makes the following contributions:
(1) We introduce MobileGeo, a novel method built upon the precision-focused efficient design paradigm. This approach resolves the accuracy-efficiency dilemma in CVGL by concentrating model complexity during training, yielding a highly accurate inference model that is lightweight enough for deployment on mobile resource-constrained devices.
(2) We introduce Hierarchical Distillation for CVGL (HD-CVGL), a novel training framework that synergistically combines inverse self-distillation, uncertainty-aware alignment, and cross-distillation to create a compact feature extractor that excels at capturing both semantic and spatial information.
(3) We propose the Multi-view Selection Refinement Module (MSRM) and provide a theoretical demonstration grounded in mutual information explaining how it enhances localization by optimally selecting and fusing multi-view information while minimizing feature matching overhead.
(4) We conduct extensive empirical evaluations on widely used benchmarks, including University-1652 and SUES-200, demonstrating that MobileGeo establishes a new state-of-the-art in both accuracy and efficiency. Additional deployment on edge devices further validate its real-time capabilities in the context of Mobile Edge Computing (MEC).
II Related work
II-A Cross-view drone Geo-localization
Cross-View Geo-Localization involves matching a query image against a geo-referenced database from another view. Early methods relied on handcrafted descriptors like SIFT [16] or global scene descriptors like HoG [17], but showed limited robustness to the extreme cross-view domain gap. The paradigm shifted with deep metric learning, where Siamese or Triplet networks [18, 19] learn a shared, view-invariant embedding space. To explicitly address geometric inconsistencies, some works integrate geometric priors, notably using a polar transform to align satellite and ground-level perspectives [20, 21] before feature comparison. While ground-to-satellite matching [22, 23] sets the foundation, the unique oblique perspective of drone introduces a more severe and spatially non-uniform multi-view set of geometric challenges.
Drone-to-Satellite Geo-Localization. This task is particularly challenging due to the multi-view oblique and low-altitude perspective of drone, which creates significant domain gaps between platforms. Following the establishment of key benchmarks [1, 13], research has progressed from improving feature robustness with attention mechanisms [24] to the now-dominant Vision Transformer (ViT) architectures. ViTs like TransGeo [7] and FSRA [25] leverage self-attention to learn powerful global representations, setting high performance standards. More recently, the field has explored more efficient and powerful backbone architectures. For instance, several works have successfully employed ConvNeXt [26] to extract highly discriminative global features, achieving a strong balance of performance and efficiency [27]. This trend highlights a focus on backbone potency for feature extraction. However, most of contemporary methods [4, 10, 14] achieve higher accuracy by incorporating sophisticated auxiliary branches or modules, this paradigm is ill-suited for resource-constrained platforms. We follow a precision-focused efficient design paradigm to empower a simple backbone. By strengthening its feature extraction capabilities without adding any inference-time modules, our model achieves competitive precision while drastically reducing the computational burden, making it ideal for deployment on platforms like drone.
II-B Multi-view Refinement
Traditional drone-based geo-localization relies on direct matching between individual query images and a reference database [28, 3, 8]. However, these approaches struggle with significant viewpoint discrepancies caused by factors such as occlusions from structures or vegetation, and the diverse perspectives captured by drone operating at varying altitudes and angles, making accurate location recognition increasingly challenging. Recent works have recognized the importance of leveraging multiple drone views to improve localization accuracy with approaches using 3D reconstruction [29, 30, 31] to represent scenes from multi-view observations and iteratively refining camera poses to align rendered views with satellite imagery. This multi-view fusion paradigm [32] demonstrates significant improvements in localization performance by exploiting the rich geometric and appearance information contained across different viewpoints.
Mutual information (MI) has emerged as a principled criterion for selecting informative views in a wide variety of complex systems. In 3D reconstruction, MI has guided next-best-view selection effectively [33]. Similarly, in multi-view clustering, recent work [34] minimizes MI between common and view-specific representations to exploit inter-view complementary information while maximizing MI between refined and original representations to preserve principal information.
Existing drone-based geo-localization methods process all available views through expensive reconstruction-based methods [15, 29], missing the opportunity for principled information-theoretic view selection. However, despite their theoretical advantages, these multi-view approaches impose substantial computational burdens that severely limit their practical drone onboard applicability. In contrast to these complex approaches with substantial computational barriers, we propose an efficient MSRM that operates as a lightweight post-processing method. Rather than constructing expensive 3D representations or performing iterative rendering optimization, our method directly aggregates selected features from multiple drone viewpoints through Mutual Information (MI).
II-C Efficient On-Device Inference for Mobile Systems
The challenge of deploying CVGL models on mobile devices is a central theme in mobile computing research [35, 36]. Knowledge Distillation (KD) has become an essential technique in transferring knowledge from large, powerful models to efficient ones, particularly in environments with constrained resources, such as Mobile Edge Computing (MEC). Several approaches [37, 38] have explored this concept in the context of edge intelligence. For example, Xu et al. [39] proposed a heterogeneous federated learning framework, driven by multi-knowledge distillation, which facilitates efficient knowledge transfer between models with different architectures. In another study, Zhang et al. [40] introduced a bidirectional knowledge distillation method, enhancing the knowledge transfer process by simultaneously distilling knowledge from both the teacher and student models. Additionally, Luo et al. [41] proposed KeepEdge, a knowledge distillation-empowered edge intelligence framework for UAV delivery. Their work demonstrates the practical application of KD in improving edge intelligence for UAVs in MEC systems.
As mentioned above, the cross model distillation from a foundation model provides general-purpose guidance, but may not fully adapt to the target task’s specific constraints. Conversely, self-distillation [42] ensures internal consistency but lacks an external source of ”expert” knowledge. Our work is situated at the intersection of these two paradigms. We propose a Hierarchical Knowledge Distillation (HD-CVGL) paradigm, enhancing robustness and generalization without requiring any external modules during inference.
III Proposed Method
In this section, we provide a comprehensive introduction to our proposed MobileGeo. As illustrated in Figure 3, we follow a precision-focused efficient design in two stages: training and inference. During the training stage, we propose Hierarchical Distillation for CVGL (HD-CVGL). Initially, we employ Fine-Grained Inverse Self-distillation (FISD) based on the internal stages of an efficient student model. Concurrently, to address the challenge of data imbalance, we introduce Uncertainty-Aware Prediction Alignment (UAPA), achieving a more robust and stable train. Subsequently, guided by a fine-tuned foundation model, we perform cross-distillation to obtain a student model with higher precision. In the inference stage, we propose a Multi-view Selection Refinement Module (MSRM) to better leverage multi-view information and further reduce the time cost of feature matching.
III-A Hierarchical Distillation for CVGL
Fine-Grained Inverse Self-distillation
In the CVGL task, a fundamental trade-off in designing deep networks exists between semantic abstraction and spatial fidelity. As shown in Figure 3 (c), the student network , composed of hierarchical stages (), transforms an input image into a sequence of feature representations , where . As the depth increases, gains semantic abstraction at the cost of losing the fine-grained spatial details present in the shallower features. In cross-view matching, these discarded low-level details often contain critical view-invariant cues (e.g., rooftop textures, landmark patterns).
To ensure the final representation is both semantically robust and perceptually detailed, we propose a novel method named fine-grained inverse self-distillation (FISD), a form of hierarchical consistency regularization. This approach inverts the conventional knowledge transfer paradigm [43]. Instead of the deep layer teaching the shallow, we compel the final student layer to retain the discriminative knowledge discovered by the shallower ”teacher” layers. This inverse knowledge transfer is motivated by two fundamental observations:
Targeting the Task-Specific Layer: It is the final layer’s feature that is ultimately used for the matching task. Consequently, this is the representation that must be refined and enriched to maximize performance.
Leveraging a Spatial-Detail-Preserving Teacher: Shallower layers serve as an authoritative teacher by preserving the fine-grained spatial information that deeper, more semantic layers progressively lose to abstraction.
We attach an auxiliary classification head to each stage’s feature map , producing logits . The core of FISD is to align the probability distribution of the final stage, , with those from all preceding stages .
Hybrid Loss Function. The training objectives of HD-CVGL in the first stage (FISD) are threefold, integrating multi-stage supervision with a focused deep metric learning objective on the final feature representation, as follows:
Multi-Level Supervision and Self-Distillation. To ensure that features at all levels of the hierarchy are semantically meaningful, we first apply a standard Cross-Entropy (CE) loss to the logits from each of the stages. This deep supervision loss is a weighted sum:
Further, to regularize the network, we employ intra-model inverse self-distillation. We first define the softened probability distribution for any stage using a temperature :
Then minimizes the Kullback-Leibler (KL) divergence between teacher distribution and the student distribution :
where denotes the KL divergence and are hyper-parameters that weight the contribution of each shallow teacher. This loss regularizes the learning of the final layer, forcing it to produce predictions consistent with the detail-oriented perspectives of the early stages.
Refining Embeddings with Symmetric Metric Learning. While the supervisory losses guide class predictions, we introduce a dedicated metric learning objective, , applied exclusively to the final-stage embeddings. The goal is to learn a shared, domain-invariant embedding space where features from drone-view and satellite-view images of the same landmark are aligned. Let a batch consist of pairs of geographically corresponding images, , where is a drone-view image and is its satellite-view counterpart. Let denote the final-stage feature embedding.
First, to enforce intra-domain class separability, we apply the triplet loss with hard sample mining strategy [12, 1]. This loss operates within each domain (drone and satellite) to ensure that embeddings of the same class are clustered together, while embeddings of different classes are pushed apart. It is applied to the final features from both domains:
Second, to achieve cross-view alignment, we employ a Symmetric InfoNCE Loss [44], a standard and powerful objective in CVGL. For a drone feature anchor , its corresponding satellite feature serves as the positive sample. All other satellite features in the batch, , act as negatives. The loss is computed symmetrically, using satellite features as anchors as well. For simplicity, let and . The Cross-view Symmetric Contrastive (CSC) loss is composed of two symmetric terms: a drone-to-satellite loss () and a satellite-to-drone loss (). The overall loss is their average:
Each directional loss is formulated as an InfoNCE loss. For a batch of pairs, the drone-to-satellite loss is defined as:
and the satellite-to-drone loss is its symmetric counterpart:
where is the cosine similarity and is a temperature parameter. This loss explicitly pulls the representations of corresponding drone-satellite pairs together while pushing them away from non-matching pairs. The total deep metric learning objective combines these components, applied to the final embeddings from both views:
This composite loss ensures that the final embedding space is both class-discriminative within each domain and robustly aligned across different views.
Uncertainty-Aware Prediction Alignment
In cross-view geo-localization, a significant challenge arises from the inherent imbalance and domain discrepancy between drone and satellite views. Specifically, for each drone-view query, only a single positive satellite sample exists within a large gallery, creating a severe data imbalance. Conventional methods often treat both domains equally, overlooking the fact that one view might be inherently more ambiguous than the other, leading to suboptimal feature alignment. As shown in Figure 4, our observation suggests an asymmetric convergence behavior, where the model may be specializing on the features of the dominant domain. This insight motivates balanced feature learning for robust cross-view matching. To address this, we introduce an uncertainty-aware prediction alignment mechanism that adaptively modulates the learning process based on the model’s confidence in each view.
Our approach begins by quantifying the predictive uncertainty of each domain. Inspire by [45, 43, 46], we employ Shannon entropy, a standard measure of uncertainty, calculated from the softmax probabilities derived from the model’s output logits. For a given logit vector over classes, the uncertainty is defined as:
here, represents the predicted probability for class , and is the corresponding logit. A higher entropy value signifies greater uncertainty and thus lower confidence in the prediction.
We then dynamically adjust the alignment process based on the relative uncertainty between the drone and satellite branches. We compute the uncertainties and for the respective logit predictions and . The core of our method is cross-modal self distillation with an adaptive temperature scaling strategy. The temperature is adjusted based on the uncertainty gap, :
where is a pre-defined base temperature and is the sigmoid function. The sigmoid function smoothly maps the unbounded uncertainty gap to a bounded scaling factor in the range (0, 1). This formulation increases the temperature when the drone-view model is more uncertain than the satellite-view model (i.e., ). This is critical for our self-distillation. When the drone view is ambiguous (e.g., due to view change or occlusions), its model is naturally uncertain. Forcing it to match the satellite’s high-confidence prediction would create a conflicting learning signal. Raising the temperature softens the target, mitigating this conflict and providing a more appropriate guidance for the uncertain student.
Finally, we use the adaptive temperature to guide the alignment between the two domains via a Kullback-Leibler (KL) divergence loss. The satellite branch acts as a ”teacher,” providing a soft target distribution for the drone ”student” branch. The alignment loss, , is formulated as:
By making the alignment process sensitive to predictive uncertainty, our method fosters a more robust and stable training process, effectively mitigating the challenges posed by data imbalance and domain-specific ambiguity.
Cross-distillation training
The second step of our hierarchical framework is Cross-Distillation Training (CDT), a process designed to transfer knowledge from a large foundation model (DINOv2-base [47]) teacher to a lightweight ConvNeXt-tiny [26] student, as illustrated in Figure 3 (a). This approach adheres to the MobileGeo paradigm by directly training the final student architecture without requiring any inference-time modifications. Critically, this teacher is not used off-the-shelf; it is first specialized for the target task through a parameter-efficient fine-tuning process on the University-1652 dataset, where only the final two Transformer blocks were made trainable. This approach creates an expert teacher that retains vast, general visual knowledge while acquiring high-level semantic understanding specific to our domain.
To ensure comprehensive knowledge transfer, we distill information at both the feature and logit levels as follows:
where denotes normalization, denote the teacher’s and student’s final stage output feature, are their respective temperature-scaled probability outputs, and is the ground-truth label. By distilling both feature semantics and classification logic from the final outputs of a domain-adapted teacher, our student model learns highly relevant representations and decision boundaries for the task.
III-B Multi-view Selection Refinement Module (MSRM)
In drone-based visual geo-localization, capturing multiple viewpoints of landmarks is essential for robust matching against satellite references. During data collection, drones systematically capture images at predetermined positions, resulting in a comprehensive multi-view representation. Let denote the set of aerial views captured at different heights and azimuth angles . While this dense sampling ensures complete coverage of landmarks, processing all views during inference poses significant computational challenges, particularly for real-time navigation applications.
To address this challenge, as shown in Figure 5, we propose the MSRM, a post-processing technique that intelligently selects an optimal subset with . The key insight is that not all views contribute equally to geo-localization accuracy: some perspectives capture more distinctive features, while others may be less informative due to occlusions or viewing angles. We formulate this view selection problem within an information-theoretic framework, maximizing the mutual information between landmark identities while ensuring spatial diversity.
Theoretical Foundation. Our approach is grounded in information theory [48, 49], where the goal is to select views that maximize the mutual information between view features and landmark labels :
where is the differential entropy of view features and is the conditional entropy given the landmark class. Direct computation of mutual information for high-dimensional features is computationally prohibitive [50]. We propose an efficient approximation based on the theoretical connection between Fisher discriminant ratio and mutual information under structured assumptions.
Proposition 1. Under the assumption that view features follow class-conditional Gaussian distributions with equal covariance, the mutual information can be lower-bounded by:
where and are the between-class and within-class variances for view .
This theoretical insight enables us to use computationally efficient statistics as proxies for mutual information while maintaining theoretical rigor.
Information-Theoretic View Importance Assessment. Given extracted multi-view features for a landmark, the MSRM quantifies each view’s information content through three complementary measures grounded in information theory. First, we approximate the marginal entropy of view features through log-variance:
This measure captures the information richness of the view, with higher entropy indicating more diverse visual patterns.
Second, we estimate the dynamic entropy through the log-range of feature activations:
This complements the variance-based entropy by capturing the span of feature activations, identifying views with strong, distinctive features.
Most importantly, we compute the Gaussian-MI approximation to measure geo-discriminability:
where the between-class variance quantifies separation between different landmarks:
and the within-class variance measures consistency within each landmark class:
here, denotes the number of samples in class , is the mean feature for class in view , is the global mean, and is the total number of samples. This formulation provides a computationally efficient estimate of mutual information while maintaining theoretical guarantees under Gaussian assumptions.
The final information score integrates these measures, where denotes min-max normalization:
Submodular Optimization for Spatial Diversity. While information scores identify informative views, optimal subset selection must balance information content with spatial coverage. We formulate this as a submodular optimization problem that jointly maximizes information and diversity.
We model each view’s spatial position as , where represents altitude and the azimuth angle. The spatial distance between views incorporates both vertical and angular separation:
where accounts for circular angles, with weights , .
Proposition 2. The objective function
is submodular, and the greedy algorithm achieves a -approximation to the optimal subset.
The greedy selection iteratively adds views that maximize the marginal gain:
with balancing information content and spatial diversity.
Information-Weighted Multi-view Aggregation. After selecting the optimal subset , we perform information-weighted aggregation that reflects each view’s contribution to the mutual information. The aggregation weights are computed using a softmax function over information scores:
where is a temperature parameter controlling the sharpness of the weighting. This exponential weighting amplifies the contribution of high-information views while maintaining differentiability. The refined representation for landmark is:
Theoretical Analysis and Guarantees. Our manual information based framework provides the following theoretical guarantees:
Approximation Quality: Under Gaussian assumptions, the approximation error is bounded by where measures deviation from Gaussianity.
Computational Efficiency: By reducing the feature matching complexity from to , our MSRM achieves a significant speedup. is the number of drone views of one landmark, this corresponds to a reduction in computational cost, rendering the approach viable for real-time applications.
The effectiveness of our approach stems from a principled connection between mutual information and discriminative learning. This connection enables efficient view selection, preserving the most informative perspectives while ensuring comprehensive spatial coverage of drone-view landmarks.
IV Experiment
IV-A Implementation Details
Datasets. We conduct extensive experiments on two prominent drone-based geo-localization benchmarks that offer complementary characteristics for comprehensive evaluation. University-1652 [1] is the first drone-based geo-localization dataset and SUES-200 [13] represents a pioneering benchmark that considers aerial photography captured by drones at different flight heights. These datasets collectively enable comprehensive evaluation of cross-view geo-localization algorithms under varying conditions—University-1652 excels in scale and multisource integration, while SUES-200 provides critical insights into altitude robustness and real-world applicability.
Experimental Setup. We train our model using a batch size of , where each batch contains different location IDs with samples per ID. This results in drone images and satellite images per batch. All images are resized to pixels for both training and testing phases. The model is trained for epochs using the SGD optimizer, initialized with a learning rate of and incorporating a 5 epoch warm-up phase following [51] to stabilize early-stage gradient dynamics. For cross-region evaluation in the unsupervised domain adaptation setting, we implement a contrastive learning fine-tuning strategy, limiting adaptation to a single epoch on the University-1652 dataset to mitigate distribution-specific overfitting. We employ the AutoAugment [52] strategy, which applies optimized combinations of augmentation operations including random cropping, translation, horizontal flipping, and color jittering. All experiments are conducted using the PyTorch framework on a single NVIDIA GeForce RTX 3090 GPU with 24GB memory.
Evaluation Protocol. Following standard practice in cross-view geo-localization [1, 7], we evaluate both retrieval accuracy and computational efficiency across all models.
Retrieval Metrics. We adopt Recall@1 to measure the fraction of queries where the correct result appears in the top position. Average Precision (AP) provides a comprehensive assessment by considering all relevant items in the ranking.
Efficiency Metrics. We report three efficiency metrics: (1) Parameters: the total number of learnable parameters in millions; (2) FLOPs: floating-point operations computed for a single forward pass; (3) Processing Speed (FPS): in the feature extraction context, the number of drone and satellite images processed per second. Note that we set the batch size to 16 to accommodate the multi-view drone data processing.
| Method | Drone Satellite | Satellite Drone | |||||
| Recall@1 | AP | Recall@1 | AP | ||||
| LPN [2] (Wang et al. 2021) | 62.39 M | 65.39 G | 218 | 77.71 | 80.80 | 90.30 | 78.78 |
| FSRA [8] (Dai et al. 2021) | 53.16 M | 98.05 G | 100 | 85.50 | 87.53 | 89.73 | 84.94 |
| MCCG [28] (Shen et al. 2023) | 56.65 M | 51.04 G | 313 | 89.40 | 91.07 | 95.01 | 89.93 |
| MuSe-Net [53] (Wang et al. 2024) | 82.90 M | 42.37 G | 405 | 74.48 | 77.83 | 88.02 | 75.10 |
| SCPNet [54] (Gao et al. 2025) | - | - | - | 79.96 | 83.04 | 87.33 | 79.87 |
| TriSA [3] (Sun et al. 2024) | 51.13 M | 43.18 G | 275 | 90.08 | 91.56 | 96.01 | 90.12 |
| Safe-Net [55] (Lin et al. 2024) | 52.67 M | 24.58 G | 282 | 86.98 | 88.85 | 91.22 | 86.06 |
| SDPL [9] (Chen et al. 2024) | 42.56 M | 69.71 G | 519 | 90.16 | 91.64 | 93.58 | 89.45 |
| SRLN [56] (Lv et al. 2024) | 193.03 M | - | - | 92.70 | 93.77 | 95.14 | 91.97 |
| Sample4Geo [14] (Deuser et al. 2023) | 87.57 M | 90.24 G | 144 | 92.65 | 93.81 | 95.14 | 91.39 |
| CCR [11] (Du et al. 2024) | 156.57 M | 160.61 G | - | 92.54 | 93.78 | 95.15 | 91.80 |
| DAC [10] (Xia et al. 2024) | 96.50 M | 90.24 G | 128 | 94.67 | 95.50 | 96.43 | 93.79 |
| MEAN [4] (Chen et al. 2025) | 36.50 M | 26.18 G | 307 | 93.55 | 94.53 | 96.01 | 92.08 |
| MobileGeo (Ours) | 28.57 M | 4.45 G | 1022 | 93.87 | 94.83 | 95.72 | 92.57 |
| MobileGeo (Ours w/ Post-process) | 28.57 M | 4.45 G | 1022 | 97.15 | 97.50 | 95.58 | 96.27 |
IV-B Comparison with State-of-the-Art Methods
We conduct a comprehensive evaluation of our proposed MobileGeo framework against a wide array of recent state-of-the-art methods. As presented in Table I, the comparison focuses on both retrieval accuracy (Recall@1, AP) and model efficiency (Parameters, FLOPs, FPS).
Superior Efficiency. Our core model, MobileGeo, exhibits high computational efficiency. With only 28.57M parameters and an exceptionally low 4.45G FLOPs, it is by far the most lightweight and computationally inexpensive model among all compared methods. To put this in perspective, compared to the recent efficient model MEAN [4], MobileGeo reduces FLOPs by a remarkable factor of 5.8 and parameters by 21.2%. This optimization directly translates to a massive 3.3 increase in FPS, reaching 1022 images/second, a critical capability for real-time deployment on resource-constrained platforms.
State-of-the-Art Accuracy. Building upon this highly efficient foundation, our full model, MobileGeo†, incorporates the MSRM as a post-processing step achieves an impressive 97.15% Recall@1 and 97.50% AP in the primary DroneSatellite retrieval task. This represents a substantial absolute improvement of 3.30% in R@1 over our efficient baseline and significantly surpasses the previous best-performing method, DAC [10], all while operating with over 20 fewer FLOPs (4.45G vs. 90.24G).
IV-C Unsupervised Domain Adaptation Results
To rigorously assess the generalization capabilities of our model, we conducted zero-shot evaluations by training on the University-1652 dataset and directly testing on the SUES-200 dataset without any fine-tuning. This challenging setting simulates real-world deployment where models must handle unseen data domains. The comparative results are detailed in Table II (DroneSatellite) and Table III (SatelliteDrone).
As demonstrated in Table II, MobileGeo exhibits exceptional generalization in the DroneSatellite task. While MEAN [4] shows strong performance at 150m, our model consistently outperforms all state-of-the-art methods at 200m, 250m, and 300m. Notably, at the 300m range, MobileGeo achieves a top-ranking R@1 score of 97.87, surpassing the next-best competitor by a significant 3.24 percentage points. It is crucial to highlight that MobileGeo utilizes only 4.45G FLOPs, compared to the 26.18G FLOPs required by MEAN.
The superior generalization of MobileGeo is more pronounced in the SatelliteDrone task, as shown in Table III. In this scenario, our model unequivocally achieves the best performance across all evaluation altitudes. For instance, it surpasses the strong baseline DAC [10] by 3.75 percentage points in R@1 at the 250m altitude.
IV-D Multi-weather drone imagery degradation Results
As shown in Figure 6, to evaluate the model’s robustness in practical scenarios, we conducted extensive experiments on the University-1652 dataset under various environmental degradations, including fog, rain, darkness, and wind. As presented in Table IV, these tests assess the model’s ability to maintain high accuracy even when the visual quality of drone imagery is severely compromised.
In the DroneSatellite retrieval task, our proposed MobileGeo demonstrates remarkable resilience and establishes a new state-of-the-art across all tested conditions. From normal weather to the most severe degradations like darkness and combined fog with rain, MobileGeo consistently achieves the highest Recall@1 and AP scores. For instance, under dark conditions, it outperforms the next-best method, MEAN [4], by a substantial margin of 5.37 percentage points in R@1 (93.27% vs. 87.90%).
In the more challenging SatelliteDrone task, MobileGeo continues to show highly competitive performance. In the most complex scenarios such as ”Fog+Rain” and ”Wind”, MobileGeo takes the lead in both metrics, achieving the best R@1 and AP scores. This proves that even when satellite images must be matched against heavily distorted drone views, our model’s feature representation remains robust and effective.
| DroneSatellite | |||||||||||
| 150m | 200m | 250m | 300m | ||||||||
| Model | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP | |||
| MCCG [28] (Shen et al. 2023) | 56.65 M | 51.04 G | 313 | 57.62 | 62.80 | 66.83 | 71.60 | 74.25 | 78.35 | 82.55 | 85.27 |
| Sample4Geo [14] (Deuser et al. 2023) | 87.57 M | 90.24 G | 144 | 70.05 | 74.93 | 80.68 | 83.90 | 87.35 | 89.72 | 90.03 | 91.91 |
| DAC[10] (Xia et al. 2024) | 96.50 M | 90.24 G | 128 | 76.65 | 80.56 | 86.45 | 89.00 | 92.95 | 94.18 | 94.53 | 95.45 |
| MEAN [4] (Chen et al. 2025) | 36.50 M | 26.18 G | 307 | 81.73 | 87.72 | 89.00 | 91.00 | 92.13 | 93.60 | 94.63 | 95.76 |
| MobileGeo (Ours) | 28.57 M | 4.45 G | 1022 | 79.85 | 83.32 | 91.10 | 92.79 | 95.57 | 96.45 | 97.87 | 98.32 |
| SatelliteDrone | |||||||||||
| 150m | 200m | 250m | 300m | ||||||||
| Model | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP | |||
| MCCG [28] (Shen et al. 2023) | 56.65 M | 51.04 G | 313 | 61.25 | 53.51 | 82.50 | 67.06 | 81.25 | 74.99 | 87.50 | 80.20 |
| Sample4Geo [14] (Deuser et al. 2023) | 87.57 M | 90.24 G | 144 | 83.75 | 73.83 | 91.25 | 83.42 | 93.75 | 89.07 | 93.75 | 90.66 |
| DAC[10] (Xia et al. 2024) | 96.50 M | 90.24 G | 128 | 87.50 | 79.87 | 96.25 | 88.98 | 95.00 | 92.81 | 96.25 | 94.00 |
| MEAN [4] (Chen et al. 2025) | 36.50 M | 26.18 G | 307 | 91.25 | 81.50 | 96.25 | 89.55 | 95.00 | 92.36 | 96.25 | 94.32 |
| MobileGeo (Ours) | 28.57 M | 4.45 G | 1022 | 92.50 | 83.81 | 97.50 | 91.75 | 98.75 | 94.59 | 97.50 | 96.04 |
| Method | DroneSatellite | SatelliteDrone | |||
| FLOPs | Recall@1 | AP | Recall@1 | AP | |
| (a) Normal | |||||
| LPN [2] | 65.39 G | 74.33 | 77.60 | 87.02 | 75.19 |
| MuSeNet [53] | 42.37 G | 74.48 | 77.83 | 88.02 | 75.10 |
| Sample4Geo[14] | 90.24 G | 90.55 | 92.18 | 95.86 | 89.86 |
| MEAN[4] | 26.18 G | 90.81 | 92.32 | 96.58 | 89.93 |
| MobileGeo (Ours) | 4.45 G | 93.87 | 94.83 | 95.72 | 92.57 |
| (b) Fog | |||||
| LPN [2] | 65.39 G | 69.31 | 72.95 | 86.16 | 71.34 |
| MuSeNet [53] | 42.37 G | 69.47 | 73.24 | 87.87 | 69.85 |
| Sample4Geo[14] | 90.24 G | 89.72 | 91.48 | 95.72 | 88.95 |
| MEAN[4] | 26.18 G | 90.97 | 92.52 | 96.00 | 89.49 |
| MobileGeo (Ours) | 4.45 G | 92.95 | 94.08 | 95.72 | 91.17 |
| (c) Rain | |||||
| LPN [2] | 65.39 G | 67.96 | 71.72 | 83.88 | 69.49 |
| MuSeNet [53] | 42.37 G | 70.55 | 74.14 | 87.73 | 71.12 |
| Sample4Geo [14] | 90.24 G | 85.89 | 88.11 | 94.44 | 85.71 |
| MEAN [4] | 26.18 G | 88.19 | 90.05 | 95.15 | 88.87 |
| MobileGeo (Ours) | 4.45 G | 91.26 | 92.61 | 94.58 | 87.22 |
| (d) Fog+Rain | |||||
| LPN [2] | 65.39 G | 64.51 | 68.52 | 84.59 | 66.28 |
| MuSeNet [53] | 42.37 G | 65.59 | 69.64 | 85.02 | 67.78 |
| Sample4Geo [14] | 90.24 G | 85.88 | 88.16 | 93.44 | 85.27 |
| MEAN [4] | 26.18 G | 86.75 | 88.84 | 93.58 | 86.91 |
| MobileGeo (Ours) | 4.45 G | 90.68 | 92.11 | 94.86 | 85.67 |
| (f) Dark | |||||
| LPN [2] | 65.39 G | 53.68 | 58.10 | 82.88 | 52.05 |
| MuSeNet [53] | 42.37 G | 53.85 | 58.49 | 80.74 | 53.01 |
| Sample4Geo[14] | 90.24 G | 87.90 | 89.87 | 96.01 | 87.06 |
| MEAN[4] | 26.18 G | 87.90 | 89.87 | 96.29 | 89.87 |
| MobileGeo (Ours) | 4.45 G | 93.27 | 94.34 | 95.44 | 89.95 |
| (g) Wind | |||||
| LPN [2] | 65.39 G | 66.46 | 70.35 | 84.14 | 67.35 |
| MuSeNet [53] | 42.37 G | 69.45 | 73.22 | 86.31 | 70.03 |
| Sample4Geo[14] | 90.24 G | 83.39 | 89.51 | 95.29 | 87.06 |
| MEAN[4] | 26.18 G | 89.27 | 91.01 | 95.44 | 86.05 |
| MobileGeo (Ours) | 4.45 G | 93.45 | 94.48 | 95.58 | 91.61 |
IV-E Anti-offset Generalization Results
In real-world drone-view geo-localization, the captured query image is often not perfectly centered or aligned with its corresponding satellite-view gallery image. This spatial misalignment can be caused by variations in camera angle or differences in viewpoint. A robust model must be able to generalize well despite such spatial offsets.
To rigorously evaluate our model’s robustness against this challenge, we adopt the evaluation protocol popularized by SDPL [9]. As shown in Figure 7, the mapping from a desired shift to the required (left, top, right, bottom) padding tuple is as follows: top-left , top-right , bottom-left , and bottom-right .
The comprehensive results are presented in Table V. The findings demonstrate the superior robustness of our proposed MobileGeo model. For example, under the most severe shift of pixels, where other models experience a significant performance collapse, MobileGeo maintains an exceptional Rank-1 accuracy of 93.86%. This represents a massive improvement of +16.92% over the second-best method, SDPL.
It is crucial to note that while the performance of all models degrades as the offset magnitude increases, MobileGeo’s degradation is significantly more graceful.
| Padding Pixel | FSRA (98.05 G) | LPN (65.39 G) | SDPL (69.71 G) | MobileGeo (4.45 G) | ||||
| Recall@1 | AP | Recall@1 | AP | Recall@1 | AP | Recall@1 | AP | |
| (0,0) | 86.41 | 88.34 | 77.99 | 80.84 | 85.19 | 87.43 | 93.87 (+7.46) | 94.83 (+6.49) |
| (+20,0) | 85.51 | 87.59 | 76.64 | 79.62 | 84.42 | 86.78 | 93.19 (+7.68) | 94.28 (+6.69) |
| (+40,0) | 82.77 | 85.30 | 72.94 | 76.36 | 82.46 | 85.13 | 90.84 (+8.07) | 92.34 (+7.04) |
| (+60,0) | 77.95 | 81.18 | 66.85 | 70.91 | 79.22 | 82.43 | 85.17 (+5.95) | 87.59 (+5.16) |
| (-20,-20) | 84.35 | 86.62 | 76.40 | 79.42 | 84.39 | 86.76 | 93.87 (+9.48) | 94.83(+8.07) |
| (-40,-40) | 78.10 | 81.24 | 70.27 | 74.03 | 81.75 | 84.55 | 93.88 (+12.13) | 94.84 (+10.29) |
| (-60,-60) | 67.73 | 71.97 | 59.56 | 64.34 | 76.94 | 80.46 | 93.86 (+16.92) | 94.82 (+14.36) |
| (+20,-20) | 84.23 | 86.55 | 76.34 | 79.37 | 84.32 | 86.72 | 93.19 (+8.87) | 94.28 (+7.56) |
| (+40,-40) | 77.90 | 81.09 | 70.36 | 74.10 | 81.62 | 84.46 | 90.85 (+9.62) | 92.35 (+8.31) |
| (+60,-60) | 67.29 | 71.62 | 59.61 | 64.42 | 76.80 | 80.38 | 85.16 (+8.36) | 87.58 (+7.20) |
| (-20,+20) | 83.46 | 85.85 | 74.74 | 77.92 | 82.95 | 85.49 | 92.72 (+9.26) | 93.88 (+8.03) |
| (-40,+40) | 74.47 | 78.05 | 65.13 | 69.32 | 77.00 | 80.46 | 86.62 (+9.62) | 88.77 (+8.31) |
| (-60,+60) | 58.05 | 63.27 | 50.19 | 55.56 | 66.87 | 71.71 | 72.36 (+5.49) | 76.25 (+4.54) |
| HD-CVGL | MSRM | DroneSatellite | SatelliteDrone | ||||
| SD | UAFA | CD | Recall@1 | AP | Recall@1 | AP | |
| ✗ | ✗ | ✗ | ✗ | 86.44 | 88.69 | 93.72 | 85.13 |
| ✓ | ✗ | ✗ | ✗ | 91.46 | 92.91 | 94.57 | 90.49 |
| ✓ | ✓ | ✗ | ✗ | 91.93 | 93.29 | 95.14 | 91.35 |
| ✓ | ✓ | ✓ | ✗ | 93.87 | 94.83 | 95.72 | 92.57 |
| ✓ | ✓ | ✓ | ✓ | 97.15 | 97.50 | 95.58 | 96.27 |
IV-F Ablation Studies
Our MobileGeo framework is architected training-time optimization and lightweight inference-time post-process. To validate the specific contributions, we conducted extensive ablation studies, with the results summarized in Table VI.
Effectiveness of the Training-Phase Framework. The core objective of our training strategy is to distill rich, view-invariant knowledge into a single, efficient network, avoiding the need for multi-branch architectures during inference. Our analysis starts with a baseline model (row 1) achieving 86.44% Recall@1. By introducing Self-Distillation (SD), performance improves significantly to 91.46%, which ensures that the model can fully leverage its own multi-level feature hierarchy for more robust representation learning. Following this, integrating the UAFA module further advances the Recall@1 to 91.93%. Finally, Cross-Distillation (CD) elevates performance to 93.87%, completing the HD-CVGL framework. This step-wise improvement validates our training-phase design, which produces a powerful feature extractor ready for efficient inference. As visualized in Figure 8 and Figure 9, our MobileGeo model learns a more discriminative feature distribution compared to the baseline. The features produced by MobileGeo exhibit smaller intra-class variance and larger inter-class distance, indicating a well-structured feature space.
Effectiveness of the Inference-Phase Module. After establishing a strong student model through our training strategy, we evaluate the contribution of the MSRM. The final row of Table VI shows the impact of applying MSRM to the descriptors generated by our fully trained model. The result is a remarkable jump in performance to 97.15% Recall@1. This +3.28% gain over the already powerful base model demonstrates that by refining the feature set at inference time, MSRM significantly enhances localization precision, fulfilling the ”Precision-Focused” aspect of our MobileGeo paradigm without altering the underlying network architecture. Figure 10 illustrates the performance of different selected views. The model achieves its peak performance with 40 views. Even at its lowest setting (20 views), our approach still significantly outperforms the baseline without the MSRM module. As illustrated by the top-5 retrieval results on the University-1652 and SUES-200 datasets (Figure 11 and Figure 12), MobileGeo demonstrates strong generalization and robustness in cross-view image matching capabilities, enabling accurate localization even in GPS-denied outdoor environments.
Edge Device Deployment. To demonstrate the practical applicability of our approach, as shown in Figure 14, we evaluate MobileGeo on two representative edge platforms: NVIDIA Jetson TX2 and AGX Orin. As shown in Figure 13, our method achieves exceptional efficiency with only 4.45 GFLOPs, representing a 95.1% reduction compared to DAC. This dramatic reduction translates directly to superior real-time performance: MobileGeo achieves 27.1 FPS on the resource-constrained TX2 (7.7 faster than DAC’s 3.5 FPS) and an impressive 251.5 FPS on AGX Orin (6.1 faster than DAC). Furthermore, even when compared to the efficient MEAN method, our approach exhibits substantial speedups. On the TX2, MobileGeo is 2.9 times faster, and on the AGX Orin, it achieves a 2.8 speedup. These results underscore the exceptional performance of MobileGeo in resource-constrained environments, highlighting its potential for real-world applications on mobile and edge devices.
As mentioned above, the efficiency gains observed are not only a result of reduced computational overhead but also reflect the method’s ability to optimize resource usage effectively, a critical factor in edge computing.
IV-G Limitations and Future Works
Despite its promising performance, MobileGeo has several limitations that guide our future work. The current model, while excelling at DroneSatellite task exhibits an asymmetric performance due to its focus on the multi-view refinement of drone features. Future efforts will explore the intra-relationships within the satellite data to build a more generalized cross-view feature space. Additionally, MobileGeo’s lightweight design faces challenges with extreme low-altitude distortions, which could be mitigated by developing adaptive, scale-aware mechanisms or pre-transformation techniques. To bridge the gap between efficient image-level retrieval and the need for precise localization, our ultimate goal is to evolve beyond static images. We plan to leverage multiview sequence data and develop an end-to-end framework that integrates coarse scene identification with fine-grained pose estimation. This would enable robust and accurate geo-localization within large-scale satellite imagery database.
To further enhance the real-time capabilities of MobileGeo, we aim to incorporate dynamic updates directly on mobile devices. By leveraging the capabilities of 5G networks and cloud-edge collaboration, we plan to optimize the model through distributed training, enabling seamless real-time geo-localization even in challenging unseen environments. This strategy will ensure that MobileGeo can not only handle large-scale datasets but also adapt quickly to changing geographical conditions, providing both efficiency and precision in mobile and drone-based applications.
V Conclusion
In this paper, we introduced MobileGeo, a mobile-friendly framework that follows a precision-focused efficient design paradigm. Our core strategy shifts computational complexity to the training stage, enabling superior performance on resource-constrained devices. We achieve this through two innovations: 1) A comprehensive Hierarchical Distillation (HD-CVGL) strategy during training, which incorporates our Uncertainty-Aware Prediction Alignment (UAPA) to robustly handle data imbalance and domain discrepancies, producing a highly discriminative yet compact student network without any inference overhead. 2) A lightweight Multi-view Selection Refinement Module (MSRM) at inference, which uses mutual information to select and fuse the most informative views, boosting accuracy while minimizing feature matching cost. Extensive experiments demonstrate that MobileGeo outperforms previous state-of-the-art methods, e.g., achieving improvements of 4.19% in AP on University-1652 while being over 5 efficient in FLOPs and 3 faster during inference. Crucially, its 251.5 FPS on AGX Orin validates its practicality for outdoor drone geo-localization in GNSS-denied environments on resource-constrained mobile and edge devices.